Given the information below, complete the chart and answer the questions that follow. Assume there are...

  1. Given the information below, complete the chart and answer the questions that follow. Assume there are no taxes, so Y = Yd. Also, assume that I, G and (X-M) are autonomous expenditures. In this problem, the MPS = .20

S

Y

C

I

G

(X-M)

TE

0

1,500

1,250

1,350

-100

5,000

10,000

15,000

20,000

25,000

what is the MPC?

At an income level of 10,000, what is the INDUCED level of consumption?

At an income level of 25,000, what is the AUTONOMOUS level of consumption?

What is the value of equilibrium income?

Using the Keynesian framework (TE/TP), plot savings and consumption (vertical axis) against income (Q or Y) on the horizontal axis
Answer #1

(a) MPS = 1 - MPC = 1 - 0.2 = 0.8

(b) Consumption function: C = a + bY

When Y = 0, C = a = 1,500

C = 1,500 + 0.8Y

When Y = 10,000,

Induced Consumption = 0.8 x 10,000 = 8,000

(c) When Y = 25,000, Autonomous consumption = 1,500 (which does not change with Y)

(d) In equilibrium, Y = TE = C + I + G + X - M

S = Y - C

S Y C I G X - M TE
-1,500 0 1,500 1,250 1,350 -100 4,000
-500 5,000 5,500 1,250 1,350 -100 8,000
500 10,000 9,500 1,250 1,350 -100 12,000
1,500 15,000 13,500 1,250 1,350 -100 16,000
2,500 20,000 17,500 1,250 1,350 -100 20,000
3,500 25,000 21,500 1,250 1,350 -100 24,000

So, equilibrium Y = 20,000

NOTE: As per Chegg Answering Policy, 1st 4 parts are answered.